OpenStudy (anonymous):

a peawingspan is sitting on top of a pillar ,which is 9m high.frm a point 27m away frm da bottom of da pillar,a snake is comming to his hole at the base of the pillar.seeing the snake the peawingspan pounces on it,if their speed are equal at wht distance from the hole is the snake caught.

OpenStudy (anonymous):

hurry up

OpenStudy (anonymous):

(27-a)^2+9^2=a^2 

OpenStudy (anonymous):

27^2-54a+a^2+9^2=a^2 27^2-54a+9^2=0 54a=27^2+9^2

OpenStudy (anonymous):

a=15 27-15=12